# Active grid turbulence anomalies through the lens of physics informed neural networks

### v1

@@ Abstract
Active grids operated with random protocols are a standard way to generate large Reynolds number turbulence in wind and water tunnels. But anomalies in the decay and third-order scaling of active-grid turbulence have been reported. We combine Laser Doppler Velocimetry and hot-wire anemometry measurements in a wind tunnel, with machine learning techniques and numerical simulations, to gain further understanding on the reasons behind these anomalies. Numerical simulations that incorporate the statistical anomalies observed in the experimental velocity field near the active grid can reproduce the experimental anomalies observed later in the decay. The results indicate that anomalies in experiments near the active grid introduce correlations in the flow that persist for long times, and result in the flow being statistically different from homogeneous and isotropic turbulence.
@@ 1 Introduction
One of the most widespread uses of active grids is to generate moderate-to-high Reynolds numbers in a wind tunnel with large values of turbulent intensity. By operating the blades in various random protocols, the lack of a characteristic time scale at the turbulence generation results in integral time and length scales that are on the order of the wind tunnel’s lateral size instead of the mesh size, as is the case with static grids [1]. This strategy allows to increase the separation of scales, reaching Reynolds numbers based on the Taylor microscale as large as $R_{\lambda}\sim 1500$ in standard wind tunnel facilities [10]. Furthermore, random protocols still generate turbulent flows that are close to homogeneous and isotropic turbulence (HIT) conditions [11] (meaning that they flow is as close to HIT as for regular static grids). Consequently, active grids have been used to study fundamental turbulence [12, 13, 14] and even two-phase flows [15, 16, 17, 18].
@@
with $u^{\prime}(x,t)$ the fluctuating streamwise velocity. While for homogeneous isotropic turbulence, within the Kolmogorov phenomenology, the value of $S_{3}(\ell)/\left(\varepsilon\ell\right)$ should be equal or lower than 0.8 [20], experiments in active grids have reported results above this number [21] for streamwise distances $x$ as large as $x/M=30$ (with $M$ the mesh size of the active grid). For larger streamwise distances ($x/M>80$), this anomaly is no longer observed [11]. Note that for passive grids, turbulence is expected to be fully developed in the range $5020<x/M<$ 50 [22, 23], whilst for active grids such distance is expected to be much shorter [24, 10]. While some of the reported anomalies are probably related to the persistence of inhomogeneities and/or anisotropy, no clear explanation has been provided to assess the anomaly in $S_{3}(\ell)$ in the near field of the wake.
@@
Another issue that arises when comparing active-grid turbulence generated with random protocols to other turbulent flows is the difficulty in defining time and length scales that would allow to compare different statistics among them. For instance, the autocorrelation function of velocity also presents anomalies, as in some cases it never crosses zero, making it difficult to define an integral length scale [25, 26]. The lack of clearly defined length scales also complicates the delineation of the production range in the near field of the turbulent flow. Moreover, the very high values of turbulence intensity near the grid (which can reach 50% or even more) make it challenging for some standard collection techniques, such as hot-wire anemometry, to properly quantify the flow. Indeed, the use of Taylor’s hypothesis in these flows presents certain limitations [27]. Furthermore, the large scale separation in these flows makes them also extremely difficult to be characterised via standard optical techniques.
@@
The present work aims at gaining further understanding about the decay of active-grid-generated turbulence using random protocols. Given the significant problems related to characterising this flow experimentally, we turn as well to numerical simulations. In practice, Laser Doppler Velocimetry (LDV) is initially used to characterise the turbulent flow at $x/M=3$. This technique results in a resolved two-dimensional (2D) map of velocity (including the average value and higher moments) in the plane perpendicular to the freestream velocity. In order to generate proper initial conditions for 3D DNSs that can resemble the near-active-grid flow in a statistical sense, we use a protocol which combines Physics-Informed Neural Networks (PINNs) [28] with a data assimilation technique known as nudging [29, 30]. The combination of the two has been shown to be successful at generating turbulence-compatible velocity fields with fixed given statistical moments [31]. This approach allows us to create a 3D box that is evolved in time using a DNS code, fully capturing and characterising the decay of turbulence. For the sake of comparison, we also tested synthetic initial conditions that correspond to HIT. Hot-wire anemometry (HWA), performed at $x/M=30$, allows us to test small-scale information (such as $S_{3}(\ell)$) in the wind tunnel experiment, aiming at verifying the anomalies previously reported in active-grid-generated turbulence, and matching the timescales between DNSs and experiments.
@@ 3.1 PINN and nudging protocol for initial conditions
As discussed above, a combination of PINNs and the nudging data-assimilation technique will be used to prepare initial conditions for the DNSs compatible with the observations near the AG. We briefly recall the main points of the used method, which was introduced and validated using synthetic data in [31]. A PINN is a neural network in which the loss term in the training is combined with physical information of a given system. For instance, a prediction may be penalised so that it satisfies a given physical law, or so that it is the solution to a given differential equation (such as the Navier-Stokes equation). In the implementation used in this work, the PINN generates predictions which are compatible with an evolution given by the incompressible Navier-Stokes equations, and whose centralised third-order moment is compatible with the observed PDF asymmetry in the wind tunnel at $x=3M$. To do this, we use a loss function
@@
Figure 3: (a) Time decay of the kinetic energy for the PINN-generated states and for HIT simulations. (b) Time evolution of the Kolmogorov wavenumber, normalised by the largest wavenumber resolved by the simulations. The vertical dotted line indicates the time where $k_{\eta}$ peaks for the PINN-generated states, which is also indicated in panel (a). (c) Evolution of the Taylor-scale Reynolds number. (d) Integral length scale $L_{0}$ as a function of time. In all panels the thick solid and dotted lines correspond to the mean value over the 10 realisations, while the shaded bands represent typical variations between the different realisations; labels for all panels correspond to the ones in panel (a).
@@ 3.2 Free decay of prepared states
Figure 4: (a) Third-order longitudinal structure function, compensated by $\varepsilon\ell$, for the HIT runs and for the PINN-generated states, at time $t/T_{0}\approx 5$. The thick solid and dotted lines correspond to the mean value over the 10 realisations, while the shaded bands represent typical variations between the different realisations. (b) Evolution of the third-order moment of the streamwise component of the velocity, normalised by the value imposed during the PINN protocol. Thick and shaded regions are as in panel (a). The thin dotted lines correspond to the time evolution in each one of the single HIT simulations. The black solid and red dotted vertical lines indicate the time at which the structure functions in panel (a) are computed. The dashed vertical line indicated the time where $k_{\eta}$ peaks for the PINN-generated states (see Fig. 3). (c) Peak value of the normalised third-order structure function along the decay. The dash-dotted vertical line indicates the equivalent time at which the hot-wire measurements in the wind tunnel are performed. (d) Ratio of r.m.s. values of the fluctuating velocity components, shown in black and red markers for PINN-generated and HIT states, respectively.
@@
The evolution of the Taylor-based Reynolds number $R_{\lambda}$ is shown in Fig. 3(c). Even when initially $R_{\lambda}$ is larger for the PINN-states than for HIT, at around $t/T_{0}\approx 7$ the values of $R_{\lambda}$ of the two types of flows converge. In Fig. 3(d) we plot the integral length scale $L_{0}(t)$, computed as in Eq. (10) with the corresponding spectrum $E(k,t)$. As with active-grids, which are characterised by a larger integral length scale than the regular, passive grids (as the former are in the order of the tunnel lateral size and the latter of the mesh size $M$), the PINN-generated states also present a higher value of $L_{0}$ when compared with the HIT states. After $k_{\eta}$ has reached its peak, the integral length scale has a dip and then the values overlap with the those of HIT.
@@
This is further supported by considering the time evolution of the maximum value of $S_{3}/(\varepsilon\ell)$, shown in Fig. 4(c). We also indicate with a vertical line the time $t/T_{0}$ that corresponds to the hot-wire measurements performed in the wind tunnel (shown in Fig. 2). While for the HIT simulations the peak value quickly drops below $4/5$, the peak for the PINN runs is, on average, larger than the expected value for longer periods of time. Even though the experimental structure function remains anomalous for larger times than in the simulations (i.e., the peak value of $S_{3}/\varepsilon\ell$ is larger than $4/5$ for longer times), imposing the initial centralised third-order moment in turbulent states has an effect on the evolution of the structure function that qualitatively shifts the HIT behaviour towards something closer to the experiment. Finally, when looking at the ratio of r.m.s. component-wise velocities, shown in Fig. 4(d), the PINN-generated states present acceptable levels of isotropy which are comparable with experimental observations in the same wind tunnel [12].
@@ 4 Conclusions
However, anomalies observed in wind tunnel experiments [21], and difficulties in reproducing these flows in simulations when only partial statistical information is available from observations [31], have raised questions on the properties of the turbulent flow generated by active grids. We have presented a case study on how to combine data from real experiments with machine learning, data assimilation, and numerical simulations, to shed light on some of these questions.

### v2

@@ Abstract
Active grids operated with random protocols are a standard way to generate large Reynolds number turbulence in wind and water tunnels. But anomalies in the decay and third-order scaling of active-grid turbulence have been reported. We combine Laser Doppler Velocimetry and hot-wire anemometry measurements in a wind tunnel, with machine learning techniques and numerical simulations, to gain further understanding on the reasons behind these anomalies. Numerical simulations that incorporate the statistical anomalies observed in the experimental velocity field near the active grid can reproduce the experimental anomalies observed later in the decay. The results indicate that anomalies in experiments near the active grid introduce correlations in the flow that can persist for long times.
@@ 1 Introduction
One of the most widespread uses of active grids is to generate moderate-to-high Reynolds numbers in a wind tunnel with large values of turbulent intensity. By operating the blades in various random protocols, the lack of a characteristic time scale at the turbulence generation results in integral time and length scales that are on the order of the wind tunnel’s lateral size instead of the mesh size, as is the case with static grids [1]. This strategy allows to increase the separation of scales, reaching Reynolds numbers based on the Taylor microscale as large as $R_{\lambda}\sim 1500$ in standard wind tunnel facilities [10]. Moreover, for variable density facilities, values as large as $R_{\lambda}\sim 5000$ have been reported [11, 12]. Furthermore, random protocols still generate turbulent flows that are close to homogeneous and isotropic turbulence (HIT) conditions [13, 14] (meaning that the flow is as close to HIT as for regular static grids). Consequently, active grids have been used to study fundamental turbulence [15, 16, 17] and even two-phase flows [18, 19, 20, 21].
@@
with $u^{\prime}(x,t)$ the fluctuating streamwise velocity. While for homogeneous isotropic turbulence, within the Kolmogorov phenomenology, the value of $S_{3}(\ell)/\left(\varepsilon\ell\right)$ should be equal or lower than 0.8 [23], experiments in active grids have reported results above this number [24] for streamwise distances $x$ as large as $x/M=30$ (with $M$ the mesh size of the active grid). For larger streamwise distances ($x/M>80$), this anomaly is no longer observed [13]. Moreover, it has been reported that at $x/M=75$ the far field already presents two-point statistics independent of $R_{\lambda}$ [14, 25]. Note that for passive grids, turbulence is expected to be fully developed in the range $5020<x/M<$ 50 [26, 27], whilst for active grids such distance is expected to be much shorter [28, 10]. As a result, statistical anomalies in active-grid-generated flows seem to last for longer streamwise distances than for their static counterparts, before reaching a universal behaviour close to the one predicted within the Richardson-Kolmogorov phenomenology. While some of the reported anomalies are probably related to the persistence of inhomogeneities and/or anisotropy, no clear explanation has been provided to assess the anomaly in $S_{3}(\ell)$ in the near field of the wake.
@@
Another issue that arises when comparing active-grid turbulence generated with random protocols to other turbulent flows is the difficulty in defining time and length scales that would allow to compare different statistics among them. For instance, the autocorrelation function of velocity also presents anomalies, as in some cases it never crosses zero, making it difficult to define an integral length scale [29, 30]. This issue arises when purely random protocols are used, implying that the forcing imposed by the grid is statistically unsteady. The lack of clearly defined length scales also complicates the delineation of the production range in the near field of the turbulent flow. Moreover, the very high values of turbulence intensity near the grid (which can reach 50% or even more) make it challenging for some standard collection techniques, such as hot-wire anemometry, to properly quantify the flow. Indeed, the use of Taylor’s hypothesis in these flows presents certain limitations [31]. Furthermore, the large scale separation in these flows makes them also extremely difficult to be characterised via standard optical techniques.
@@
The present work aims at gaining further understanding about the decay of active-grid-generated turbulence using random protocols. Given the significant problems related to characterising this flow experimentally, we turn as well to numerical simulations. In practice, Laser Doppler Velocimetry (LDV) is initially used to characterise the turbulent flow at $x/M=3$. This technique results in a resolved two-dimensional (2D) map of velocity (including the average value and higher moments) in the plane perpendicular to the freestream velocity. Direct numerical simulations (DNSs) are then used to understand in more detail the flow physics, as they provide access to the full velocity field, with all dynamical scales properly resolved, and hence they allow computation of gradients and correlations. Note that turbulence in wind tunnels is often compared with simulations of homogeneous and isotropic turbulence. In particular, the flow at different distances from the grid is comparable to the numerical evolution of a flow in the absence of forcing (i.e., freely decaying), where the initial condition is usually consistent with a fully developed turbulent state. In our case, in order to generate proper initial conditions for 3D DNSs that can resemble the near-active-grid flow in a statistical sense, we use a protocol which combines Physics-Informed Neural Networks (PINNs) [32] with a data assimilation technique known as nudging [33, 34]. The combination of the two has been shown to be successful at generating turbulence-compatible velocity fields with fixed given statistical moments [35]. This approach allows us to create a 3D box that is evolved in time using the DNS code, fully capturing and characterising the decay of turbulence. For the sake of comparison, we also tested synthetic initial conditions that correspond to HIT. Hot-wire anemometry (HWA), performed at $x/M=30$, finally allows us to test small-scale information (such as $S_{3}(\ell)$) in the wind tunnel experiment, aiming at verifying the anomalies previously reported in active-grid-generated turbulence, and matching the timescales between DNSs and experiments. In summary, the methodology used in this study is the following: (1) We use LDV measurements to characterize the flow near the active grid. (2) We used PINNs and nudging to assimilate the statistical information of these measurements in DNSs. (3) We compare the DNSs time evolution with HWA wind tunnel measurements downstream.
@@ 2.2 Fully developed turbulence region: Hot-Wire measurements
Figure 3: (a) Time decay of the kinetic energy for the PINN-generated states and for HIT simulations. (b) Time evolution of the Kolmogorov wavenumber, normalised by the largest wavenumber resolved by the simulations. The vertical dotted line indicates the time where $k_{\eta}$ peaks for the PINN-generated states, which is also indicated in panel (a). (c) Initial kinetic energy spectrum $E(k)$ of the PINN-generated states and of the HIT simulations. (d) Kinetic energy spectrum of the PINN-states at three successive times. In all panels the thick solid and dotted lines, and the circular and triangular markers, correspond to the mean value over the 10 realisations, while the shaded bands represent typical variations between the different realisations.
@@ 3.1 PINN and nudging protocol for initial conditions
As discussed above, a combination of PINNs and the nudging data-assimilation technique will be used to prepare initial conditions for the DNSs compatible with the observations near the AG. We briefly recall the main points of the used method, which was introduced and validated using synthetic data in [35]. A PINN is a neural network in which the loss term in the training is combined with physical information of a given system. For instance, a prediction may be penalised so that it satisfies a given physical law, or so that it is the solution to a given differential equation (such as the Navier-Stokes equation). In the implementation used in this work the PINN generates predictions which are compatible with an evolution given by the incompressible Navier-Stokes equations, and whose centralised third-order moment is compatible with the observed PDF asymmetry in the wind tunnel at $x=3M$. To do this we use a loss function
@@
Figure 4: (a) Evolution of the Taylor-scale Reynolds number. (b) Integral length scale $L_{0}$ as a function of time. In both panels the thick solid and dotted lines correspond to the mean value over the 10 realisations, while the shaded bands represent typical variations between the different realisations; labels for panel (b) are the same as panel (a).
@@ 3.2 Free decay of prepared states
Figure 5: (a) Third-order longitudinal structure function, compensated by $\varepsilon\ell$, for the HIT runs and for the PINN-generated states, at time $t/T_{0}\approx 5$. The thick solid and dotted lines correspond to the mean value over the 10 realisations, while the shaded bands represent typical variations between the different realisations. (b) Evolution of the third-order moment of the streamwise component of the velocity, normalised by the value imposed during the PINN protocol. Thick and shaded regions are as in panel (a). The thin dotted lines correspond to the time evolution in each one of the single HIT simulations. The black solid and red dotted vertical lines indicate the time at which the structure functions in panel (a) are computed. The dashed vertical line indicated the time where $xk_{\eta}$ peaks for the PINN-generated states (see Fig. 3). (c) Peak value of the normalised third-order structure function along the decay. The dash-dotted vertical line indicates the equivalent time at which the hot-wire measurements in the wind tunnel are performed. (d) Ratio of r.m.s. values of the fluctuating velocity components, shown in black and red markers for PINN-generated and HIT states, respectively.
@@
We can gain further insight in this respect by comparing the kinetic energy spectrum $E(k)$ of the PINN-generated states and of the HIT simulations. The average spectra over all realisations at $t=0$ for both flow types are shown in Fig. 3(c). Both flows display more than one decade of power-law like behaviour with an exponent close to $-5/3$, compatible with fully-developed turbulent conditions. The PINN spectrum displays a slight dip close to $k\lambda\approx 1.5$, which stems from the upper bound of the Fourier band-pass filter used in the nudging stage (Eq. (8)). This dip quickly disappears as the flow freely decays, as it can be seen in panel (d) in Fig. 3, where we show three successive $E(k)$ curves. At $t/T_{0}=0.68$ the amplitude of the dip has already decreased significantly, and at $t/T_{0}=1.36$ the PINN spectrum is almost indistinguishable from that of HIT. Note that this time corresponds to the early stages of the decay, before the peak of $k_{\eta}$ occurs (indicated in panel (b) in Fig. 3). Hence, throughout the most part of the evolution, the PINN and HIT states are similar from a spectral viewpoint.
The evolution of the Taylor-based Reynolds number $R_{\lambda}$ is shown in Fig. 4(a). Even when initially $R_{\lambda}$ is larger for the PINN-states than for HIT, at around $t/T_{0}\approx 7$ the values of $R_{\lambda}$ of the two types of flows converge. In Fig. 4(b) we plot the integral length scale $L_{0}(t)$, computed as in Eq. (10) with the corresponding spectrum $E(k,t)$. As with active-grids, which are characterised by a larger integral length scale than the regular, passive grids (as the former are in the order of the tunnel lateral size and the latter of the mesh size $M$), the PINN-generated states also present a higher value of $L_{0}$ when compared with the HIT states. After $k_{\eta}$ has reached its peak, the integral length scale has a dip and then the values overlap with the those of HIT.
@@
This is further supported by considering the time evolution of the maximum value of $S_{3}/(\varepsilon\ell)$, shown in Fig. 5(c). We also indicate with a vertical line the time $t/T_{0}$ that corresponds to the hot-wire measurements performed in the wind tunnel (shown in Fig. 2). While for the HIT simulations the peak value quickly drops below $4/5$, the peak for the PINN runs is, on average, larger than the expected value for longer periods of time. Even though the experimental structure function remains anomalous for larger times than in the simulations (i.e., the peak value of $S_{3}/\varepsilon\ell$ is larger than $4/5$ for longer times), imposing the initial centralised third-order moment in turbulent states has an effect on the evolution of the structure function that qualitatively shifts the HIT behaviour towards something closer to the experiment. We note however that $R_{\lambda}$ in simulations is smaller than in experiments. As we have shown for the experiments in Fig. 2(b), the anomalous behaviour in the amplitude of $S_{3}$ depends on the value of $R_{\lambda}$, so it would be interesting to evaluate how an increase in the Reynolds number affects the numerical results (a detailed study of Reynolds effects is out of the scope of this work, see also [41]). However, it should be kept in mind that the differences in the simulations with respect to the experiment is not just in the value of $R_{\lambda}$, but also in the fact that we only impose the centralised third order moment of $u^{\prime}$. In the experiments, higher order moments of $u^{\prime}$, as well as cross-correlations between the different velocity components, may also be partially responsible for the anomalous behaviour observed downstream. Finally, when looking at the ratio of r.m.s. component-wise velocities, shown in Fig. 5(d), the PINN-generated states present acceptable levels of isotropy which are comparable with experimental observations in the same wind tunnel [15].
@@ 4 Conclusions
However, anomalies observed in the near field of the grid in wind tunnel experiments [24], and difficulties in reproducing these flows in simulations when only partial statistical information is available from observations [35], have raised questions on the properties of the turbulent flow generated by active grids when using certain protocols. We have presented a case study on how to combine data from real experiments with machine learning, data assimilation, and numerical simulations, to shed light on some of these questions.
@@
Several studies report that far from the grid (for this flow, 70 meshes downstream or more), two-point statistics, isotropy and homogeneity reach universal behaviour  [14, 13, 42]. Considering this distance, anomalies detected in the active grid thus persist significantly farther downstream when compared to those present in static grid flows. Our results confirm that, depending on the flow Reynolds number, anomalies are still present as far as 40 grid meshes downstream. This is a range relevant for several wind tunnel studies, and future studies and applications should have particular care when analysing these flows. Moreover, the methodology presented here can also be used to characterize these flows, and to find ways to ameliorate or control these effects, as active grids have a central role when trying to reach large Reynolds numbers in experiments.